OpenStudy (anonymous):

MEDAL AND FAN FOR ANSWER! Julia pays $56.70 to fill her car with 15 gallons of gas. Which equation can be used to find c, the cost per gallon?

OpenStudy (anonymous):

c=56.70/15

OpenStudy (anonymous):

thats not one of my choices.. @andypie1010

OpenStudy (anonymous):

a.c/15=56.70 b.56.70c=15 c. 56.70+c=20 d.15c=56.70

OpenStudy (anonymous):

Then it's D. Thanks for posting your answer

OpenStudy (eric_d):

Yea

OpenStudy (eric_d):